OpenStudy (princesssleelee):

What is the value of 373 using the identity (x − y)3 = x3 − 3x2y + 3xy2 − y3? Hint: 373 = (40 − 3)3; therefore, x = 40 and y = 3.

OpenStudy (frozenghost127):

hello baby

OpenStudy (thebrainzzzzz):

Lol

OpenStudy (anonymous):

\[37^3=\left( 40-3 \right)^3=40^3-3*40^2*3+3*40*3^2-3^3=?\]

OpenStudy (princesssleelee):

im not sure @surjithayer

OpenStudy (anonymous):

where is doubt? \[\left( x-y \right)^3=x^3-y^3-3xy \left( x-y \right)=x^3-y^3-3x^2y+3xy^2\]